Find the first and second derivatives of the following and simplify whenever possible:



x = 9t ^ 2 - 1 : y = 3t+1

1
Expert's answer
2021-12-19T14:27:27-0500

Let us find the first and second derivative of the following:


"x = 9t ^ 2 - 1,\ \ y = 3t+1."


It follows that "y_x'=\frac{y_t'}{x_t'}=\frac{3}{18t}=\frac{1}{6t}."


Then "y_{x^2}''=\frac{(\frac{1}{6t})_t'}{x_t'}=\frac{-\frac{1}{6t^2}}{18t}=-\frac{1}{108t^3}."